OpenStudy (anonymous):

According to the Rational Root Theorem, which could be a factor of the polynomial f(x) = 60x4 + 86x3 – 46x2 – 43x + 8? A. x – 6 B. 5x – 8 C. 6x – 1 D. 8x + 5

OpenStudy (ranga):

Find the factors of the constant: 8. Find the factors of the leading coefficient: 60 Possible roots are: \[ ±\frac{\text{Factors of 8}}{\text{Factors of 60}} \]

OpenStudy (anonymous):

Which equation can pair with x – y = –2 to create a consistent and dependent system? A. 6x + 2y = 15 B. –3x + 3y = 6 C. –8x – 3y = 2 D. 4x – 4y = 6

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

So it's b?